OpenStudy (anonymous):

Question on Solving Quadratic Equations by Factoring... I really need help on this ._. .. I get that I have to check and make sure they end up to be 0.. but I get confused as to what to do with fractions. Here's an example.

OpenStudy (anonymous):

\[(a + 2/3 )( 5a - 1/2) = 0\]

OpenStudy (anonymous):

@Comm.Dan u should help...

OpenStudy (anonymous):

I got a= -2/3 on one. And the next 5a=1/2 .. See I get confused with fractions and dividing them in order to get the variable by itself.

OpenStudy (anonymous):

I haven't done any of those operations in a while..

OpenStudy (comm.dan):

(x+2/3) (x-1/2) is the factorization

OpenStudy (comm.dan):

Then subtract 2/3 from the left side and add 1/2 to the right side

OpenStudy (anonymous):

Just saying ahead of time here.. I did it with a calculator, and got 2/3 ; -2 as my answers, as both turn out to equal 0. a=0 for both when replaced and checked. But Im taking a test on this tomorrow so I needa have it figured out before I get to testing w/o a calculator..

OpenStudy (anonymous):

It's just minimal things such as fractions that always get me..

OpenStudy (comm.dan):

Do u need anymore help? @ryuu

OpenStudy (anonymous):

\[(a+2/3) --> a= - 2/3 ; (5a - 1/2) --> 5a= 1/2 \]

OpenStudy (anonymous):

Yeah, see that's where I get confused, with the fractions. ._.

OpenStudy (anonymous):

And as how to deal with them when there's a fraction with the variable and I need to get rid of them to the other side.

OpenStudy (anonymous):

ah.. no nevermind, thanks..